Show that (C[a, b], ||f||2) is a norm
1
Expert's answer
2020-12-15T13:37:43-0500

First of all, for "f\in\mathcal{C}([a;b])" we see that

"||f||_2 = (\int_a^b |f(x)|^2 dx)^{1/2}"

is well defined and is in "[0;+\infty)" .

Now let's verify the norm properties:

  1. Separation, "||f||_2 = 0 \leftrightarrow f=0" . If "f=0, ||f||_2=0" is obvious. Now suppose "f\neq 0, |f|^2 \neq 0" is a continuous positive function and thus "\int_a^b |f(x)|^2 dx \neq 0", therefore "||f||_2\neq 0" .
  2. Linearity, "||\alpha f||_2 = |\alpha| \cdot||f||_2" . It follows from linearity of an integral : "(\int_a^b |\alpha f|^2(x) dx)^{1/2} = (|\alpha|^2 \int_a^b |f(x)|^2 dx)^{1/2}=|\alpha|\cdot||f||_2"
  3. Triangle inequality, "||f+g||_2 \leq ||f||_2 + ||g||_2" .

"\int_a^b|f+g|^2(x) dx \leq \int_a^b|f|^2(x) dx +\int_a^b|g|^2(x)dx + 2 \int_a^b |fg| (x) dx"

We have "(\int_a^b |fg|(x) dx)^2 \leq (\int_a^b |f(x)|^2 dx) \cdot (\int_a^b |g(x)|^2 dx)", as

"\frac{|fg|}{(\int_a^b |f(x)|^2 dx)^{1/2}\cdot(\int_a^b |g(x)|^2 dx)^{1/2}} \leq \frac{\frac{|f|^2}{\int_a^b |f(x)|^2 dx}+\frac{|g|^2}{\int_a^b |g(x)|^2 dx}}{2}" (by applying "2|ab|\leq a^2+b^2" to "\frac{|f|}{(\int_a^b |f(x)|^2 dx)^{1/2}}" and "\frac{|g|}{(\int_a^b |g(x)|^2 dx)^{1/2}}" ) and thus by integrating on "[a;b]" we find :

"\frac{\int_a^b|fg| (x)dx}{(\int_a^b |f(x)|^2dx)^{1/2}\cdot(\int_a^b |g(x)|^2 dx)^{1/2}} \leq 1"

Therefore "(\int_a^b |fg|(x) dx)^2 \leq (\int_a^b |f|^2 dx) \cdot (\int_a^b |g|^2 dx)" .

And we conclude :

"||f+g||_2^2 = \int_a^b|f+g|^2 \leq \int_a^b|f|^2 +\int_a^b |g|^2 +2\int_a^b |fg|"

"||f+g||_2^2 \leq \int_a^b|f|^2 +\int_a^b|g|^2 +2(\int_a^b|f|^2)^{1/2}(\int_a^b|g|^2)^{1/2}"

"||f+g||_2^2 \leq (||f||_2+||g||_2)^2"

And thus "||f+g||_2 \leq ||f||_2 + ||g||_2" .